OpenStudy (scienceteen8):

Describe how you would use the Distributive Property to simplify (39 × 5).

OpenStudy (kendricklamar2014):

Need help?

OpenStudy (scienceteen8):

yes

OpenStudy (kendricklamar2014):

What have u done so far?

OpenStudy (scienceteen8):

i actually already did this question but i have to redo it because t was wrong i can attach a photo for you to see what i put

OpenStudy (kendricklamar2014):

Ok

OpenStudy (kendricklamar2014):

Here is how I would solve it: \[39 + 5 = 44 - 4 =40x5= 200 - 5 = 195\]

OpenStudy (scienceteen8):

is that all you need to do?

OpenStudy (kendricklamar2014):

Pretty much :)

OpenStudy (scienceteen8):

ok thank you

OpenStudy (kendricklamar2014):

Your Welcome :)